a company wants to manufature gift boxes in the shape of rectangular prisms .each gift box will have a volume of 84 cubic inches the base of the rectangular prism shold be twice as long as it is wide.what dimensions should the company choose for the gift boxes in order to minimize the surface area of each box.
Found 2 solutions by solver91311, htmentor:
Answer by solver91311(24713)


Width of base:

Length of base:

Height of prism:

Volume of prism:

Surface Area:

But then:

Hence the surface area as a function of the width is

And then the first derivative of the surface area function is:



Set the derivative equal to zero:



Solve for , then calculate and


John

My calculator said it, I believe it, that settles it


Answer by htmentor(1343)

The volume of the rectangular prism is V = l*w*h = 84 in^3
The area is A = 2(l*w + l*h + w*h)
The base l is equal to twice the width: l = 2w
Thus V = 2*w^2*h and A = 2(w^2 + 2wh + wh) = 2w(w + 3h)
Expressing h in terms of w gives h = V/(2*w^2)
Thus A = 2w(w + 3V/(2*w^2))
The area will be minimized when dA/dw = 0
0 = 4w - 3V/w^2 -> w = (3V/4)^(1/3)
Substituting the value for V gives w = 3.979
Therefore l = 7.958 and h = 84/(2*3.979^2) = 2.653